OpenStudy (anonymous):

How can i prove: cos x/sin^2 x-cos x cot^2 x = cot x sin x

OpenStudy (aravindg):

sin2x=.....

OpenStudy (anonymous):

Sorry i ment sin^2(x)

OpenStudy (aravindg):

Then cot 2x=.....

OpenStudy (aravindg):

:/

OpenStudy (fibonaccichick666):

\[\frac{cos x}{sin^2 x-cos x} cot^2 x = cot x sin x\] Is this the question?

OpenStudy (anonymous):

The minus cos x is not under the division bar

OpenStudy (fibonaccichick666):

k \[\frac{cosx}{sin^2x}−cosxcot^2x=cotxsinx\] Correct?

OpenStudy (anonymous):

Yes.

OpenStudy (fibonaccichick666):

ok @AravindG it's yours, and can you help me, I'm out of patience and ideas? http://openstudy.com/study#/updates/532bb32ce4b01b0501029215

OpenStudy (aravindg):

@Lizz111 Did you try taking LCM?

OpenStudy (anonymous):

Im not sure how to do that.

OpenStudy (fibonaccichick666):

(Make the left side into one big fraction) ie. 1/3 +1/4=(1+1)/12

OpenStudy (aravindg):

(a/b)+(c/d)=(ad+bc)/bd

OpenStudy (anonymous):

Develop the left side. Replace cot^2x by (cos^2 x)/(sin^2 x) cos x/sin^2 x - [cos x*(cos^2 x)]/sin ^2 x = = [cos x(1 - cos^2 x)]/sin^2 x = (cos x*sin^2 x)/sin^2 x = cos x The question might be wrong?

OpenStudy (anonymous):

I dont believe the question is wrong, but im not sure. After i find the LCM, what do i do from there?

OpenStudy (fibonaccichick666):

first, write out your steps. Even the teeniest tiniest

OpenStudy (fibonaccichick666):

I haven't actually done it yet, but we can see if it's right or wrong

OpenStudy (anonymous):

Develop the right side: cot x*sin x = (cos x/sin x)*sin x = cos x Both the left and right side are equal to cos x, then the equation is proven.

OpenStudy (anonymous):

Can you explain this step. Where did the (1- cos^2 x) come from. I know its equal to sin^2 x. cos x/sin^2 x - [cos x*(cos^2 x)]/sin ^2 x = [cos x(1 - cos^2 x)]/sin^2 x

OpenStudy (fibonaccichick666):

You remember \(sin^2x+cos^x=1\)?

OpenStudy (anonymous):

Yes.

OpenStudy (fibonaccichick666):

so that's how we do the sub, but first we have to set it up. We have: \[\frac{cos x - cos x*(cos^2 x)}{sin ^2 x} = \frac{[cos x(1 - cos^2 x)]}{sin^2 x}\] Do you see?

OpenStudy (anonymous):

I sorta understand. I guess my algebra skills are failing me and not my precalc. But i dont understand what is allowing you to sub that in?

OpenStudy (fibonaccichick666):

oh, just that they are equal. That's it just like we can sub secx for 1/cosx

OpenStudy (fibonaccichick666):

so we pulled out a cosx from each term, then we happened to have something useful

OpenStudy (anonymous):

Wait so did you multiply the cos x and the cos ^2 x together then factor out the cos x?

OpenStudy (fibonaccichick666):

yea, sort of

OpenStudy (anonymous):

So you got cos x - cos^3 x then factored out the cos x ?

OpenStudy (fibonaccichick666):

yea

OpenStudy (anonymous):

Okay. I understand. Thank you very much! I really appreciate it!

OpenStudy (fibonaccichick666):

np, so if you want, can you step by step, justifying as you go, tell me the proof?